solve by first eliminating the decimal numbers.
0.5x + 3.4 = 0.2x + 12.4

Found 2 solutions by Earlsdon, rapaljer:
Answer by Earlsdon(6294)

You can eliminate the decimals by multiplying everything by 10.
10(0.5x + 3.4 = 0.2x + 12.4) = 5x + 34 = 2x + 124
Now subtract 2x from both sides. 3x + 34 = 124
Subtract 34 from both sides. 3x = 90
Divide both sides by 3. x = 30
Check:
0.5(30) + 3.4 = 0.2(30) + 12.4
15 + 3.4 = 6 + 12.4
18.4 = 18.4 Solution is good.

Answer by rapaljer(4671)

First multiply each side of the equation by 10 to clear the decimals from the problem.

5x + 34 = 2x + 124

Add -2x to each side to get all variables on the left side:
5x - 2x + 34 = 2x - 2x + 124
3x + 34 = 124

Subtract 34 from each side:
3x + 34 - 34 = 124 - 34
3x = 90

Divide each side by 3
x = 30